# Kernel-based optimization of measurement operators for quantum reservoir computers

### v1

@@ Abstract
Finding optimal measurement operators is crucial for the performance of quantum reservoir computers (QRCs), since they employ a fixed quantum feature map. We formulate the training of both stateless (quantum extreme learning machines, QELMs) and stateful (memory dependent) QRCs in the framework of kernel ridge regression. This approach renders an optimal measurement operator that minimizes prediction error for a given reservoir and training dataset. For large qubit numbers, this method is more efficient than the conventional training of QRCs. We discuss efficiency and practical implementation strategies, including Pauli basis decomposition and operator diagonalization, to adapt the optimal observable to hardware constraints. Numerical experiments on image classification and time series prediction tasks demonstrate the effectiveness of this approach, which can also be applied to other quantum ML models.
@@ I Introduction
Since quantum systems are inherently disturbed by measurements, finding efficient and optimal ways to extract information is a crucial part of any quantum ML model, and numerous studies have addressed this problem [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. For variational QML models, measurement operators are often chosen based on hardware constraints or heuristics. Due to the training of the internal layers, this usually does not pose a severe problem for the accuracy of the model. However, in the case of QRCs, training happens only at the readout and thus depends crucially on the information extracted by the measurement operator. Finding optimal measurement operators is thus a key challenge in QRC design [14, 15, 16, 17, 18].
@@
In the present work, we discuss the problem of determining an optimal measurement operator for QRCs using kernel-based methods. QRCs have been introduced as a hardware efficient and noise-resilient quantum ML approach [31, 32, 33, 34], and have been successfully applied in tasks such as image classification [35, 36] and time series prediction [31, 15, 37]. We demonstrate that using an optimal measurement operator enables systematic improvements of QRC performance in these tasks. Besides considering QELMs, which employ a memoryless quantum feature map and are thus close to untrained conventional quantum ML models  [^1] By “conventional”, we mean quantum ML models that have trainable internal layers instead of only a final trainable classical regression layer., we show that our method can also be applied to recurrent QRCs with internal memory. The latter systems are typically used for time series processing [31, 32] and we explore their quantum kernel representation in detail in the present work.
@@ II.1 Model
We consider a $N$-qubit model, with dimension $D=2^{N}$ of its Hilbert space $\mathcal{H}$, described by the feature map
@@
$f(\mathbf{x})=\sum_{k=1}^{m}w_{k}\mathrm{tr}[M_{k}\rho(\mathbf{x})]$ (2)
@@
Importantly, for the present purpose, the feature map (1) is assumed to be stateless in the sense that the state $\rho$ is always fully determined by the input vector $\mathbf{x}$. Such a model is naturally realized by a quantum extreme learning machine (QELM) [32, 33, 35]. This seems at first precluding QRCs with internal memory, as commonly used for time series processing [31, 32, 33], because their current state $\rho(t)$ would not only depend on the present input, but also on the inputs preceding it. To address this issue, let $\{\mathbf{z}_{0},\mathbf{z}_{1},\ldots,\mathbf{z}_{T}\}$ be a time series ($T+1$ samples in $\mathbb{R}^{d^{\prime}}$) and assume that the QRC has a memory of length $L$, which is finite due to the echo state property of a correctly configured QRC. Then, an effective input $x_{t}$ at time $t\geq L$ can be defined as
@@
This quantity can be understood as an ordered set of $L+1$ samples $\mathbf{z}_{t}$ such that processing these samples through the QRC is sufficient to obtain the unique state $\rho(t)$ associated with $x_{t}$. Accordingly, the QRC output is fully determined by any of the P=TLP=T-L unordered effective inputs $\{x_{t}\}_{t=L}^{T-1}$ and can thus be considered stateless in this setting  [^2] In practice, an initial washout/equilibration time for the QRC must be considered, which, however, does not affect the present theoretical framework.. Prescriptions like (3) to define truncated input histories are also used in kernelization of classical reservoir computing models [29, 30].
@@ II.2 Measurement operator optimization
Trainable object $\mathbf{w}\in\mathbb{R}^{M}$ (or $\mathbb{C}^{M}$) $M\in\mathcal{M}(\mathcal{H})$ $\boldsymbol{\alpha}\in\mathbb{R}^{P}$
@@
Table 2: Three training formulations for a QRC. The training objective and solution are specialized here to the case of least-squares loss. For other kinds of losses, usually no analytic expressions for the weights $\mathbf{w}^{*}$ exist.
@@
with the same loss function $\mathcal{L}$ as in Eq. 4. In practice, this optimization over $\mathcal{M}(\mathcal{H})$ can be achieved by using in (4) as observables $\{M_{k}\}$ a complete $m=D^{2}$-dimensional basis of $\mathcal{M}(\mathcal{H})$ (e.g., the Pauli basis). The associated optimal weights $w_{k}^{*}$ are given formally by the same expression as in Eq. 5, and yield the optimal observable
@@
The above unconstrained optimization procedure becomes especially natural in its dual formulation, resting on the connection between supervised quantum ML models and quantum kernels [19]. Identifying the unconstrained feature map as $\phi(\mathbf{x})=\rho(\mathbf{x})$ allows one to define the kernel $K(\mathbf{x},\mathbf{x}^{\prime})=\mathrm{tr}(\rho(\mathbf{x})\rho(\mathbf{x}^{\prime}))$, which spans a reproducing kernel Hilbert space (RKHS) [41]. The representer theorem implies that the optimizer $M^{*}$ (optimal measurement operator) of the problem (6) is of the form (see also Appendix A)
$M^{*}=\sum_{j=1}^{P}\alpha_{j}^{*}\rho(\mathbf{x}_{j})$ (8)
@@
For the least-squares loss, the optimal coefficients are available in closed form as
$\boldsymbol{\alpha}^{*}=(\mathbf{K}+\lambda\mathbbm{1})^{-1}\mathbf{y},\qquad K_{ij}=\mathrm{tr}(\rho(\mathbf{x}_{i})\rho(\mathbf{x}_{j}))$ (10)
For other loss functions, the optimal coefficients must typically be computed numerically. Independently of the loss function $\mathcal{L}$, an optimally trained QELM [Eq. 2] admits the kernel form
@@
The generalization to multi-dimensional targets values $\mathbf{y}_{i}\in\mathbb{R}^{C}$ ($i=1,\ldots,P$) is straightforward and requires a separate kernel construction like in Eq. 11 for each class (this includes multi-class classification in a one-vs-rest scheme).
When the measurement operators $\{M_{k}\}$ form a complete basis (or at least span the relevant subspace containing $M^{*}$ [16]), primal and dual formulations are equivalent and Eqs. 7 and 8 describe the same object. Expanding the optimal observable $M^{*}$ in that basis leads to a relation between the coefficients $\mathbf{w}^{*}$ and $\boldsymbol{\alpha}^{*}$ via [cf. Eqs. 49 and 61]
$w_{k}^{*}=\sum_{j=1}^{P}\alpha_{j}^{*}\mathrm{tr}(M_{k}\rho(\mathbf{x}_{j})),$ (12)
The optimization formulations are summarized in Table 2.
@@
(i) Neither the primal nor the dual optimization methods assume uncorrelated samples. Both are therefore compatible with the sliding window construction (3). (ii) Global conjugation $\rho(\mathbf{x})\mapsto U\,\rho(\mathbf{x})\,U^{\dagger}$ leaves the expression (11) unchanged and only transforms the optimal observable as $M^{*}\mapsto U^{\dagger}M^{*}U$. (iii) The constrained primal formulation can be “dualized”, too, by using the kernel induced by the constrained feature map (see Table 2). While this does not render a lower training error, it may offer computational benefits.
@@ II.3.3 Diagonalization
The prediction is the weighted average of the observed bitstrings using the eigenvalues $\{\lambda_{\mathbf{b}}\}$. Thus, the optimal observable can be implemented by a single basis rotation $V^{\dagger}$ followed by a standard computational-basis measurement. This is equivalent to measuring all $2^{N}-1$ distinct Pauli-$Z$ strings $\{Z_{\mathbf{s}}\}_{\mathbf{s}\in\{0,1\}^{N}}$, since (see Appendix C)
@@
The unitary $V$ may be treated as part of the readout (applied once just before measurement) or absorbed into the reservoir by redefining the last layer as $U_{R}\mapsto V^{\dagger}U_{R}$. This global conjugation leaves the kernel $K(\mathbf{x},\mathbf{x}^{\prime})$ and the trained predictor unchanged (see Section II.2) and only affects the hardware realization of the optimal measurement. While, strictly, using the above pre-rotated $Z$-basis measurements are sufficient and equivalent to measuring the exact $M^{*}$, we will include in our numerical experiments [Section IV] also the Pauli observable decomposition [Eq. 13], as both approaches can lead to slight differences in numerical stability.
@@ III.1 QRCs and QELMs
The standard (stateful) QRC model introduced in [31, 32] consists of a set of coherently evolving (memory/ancilla) qubits ($N_{A}$) periodically coupled to a set of input qubits ($N_{I}$). At time $t$, a classical input $\mathbf{z}_{t}\in\mathbb{R}^{d^{\prime}}$ is encoded into an input state with density matrix $\eta_{I}(\mathbf{z}_{t})$. The joint state with the coherent ancilla state $\eta_{M}(t)$ is formed, $\eta(t)=\eta_{I}(\mathbf{z}_{t})\otimes\eta_{M}(t)$, and time-evolved by a unitary $U$, giving
$\rho(t)\equiv\eta(t+1)=U\eta(t)U^{\dagger}.$ (17)
This $\rho(t)$ corresponds to the feature map $\rho(\mathbf{x})$ in Eq. 1, after imposing Eq. 3. Subsequently, the input is reset and discarded, and the memory alone is carried to the next step:
$\eta_{M}(t+1)=\mathrm{tr}_{I}\!\left[\,U\bigl(\eta_{I}(\mathbf{z}_{t})\otimes\eta_{M}(t)\bigr)U^{\dagger}\right],$ (18)
where $\mathrm{tr}_{I}$ denotes partial trace over the input register $I$.
We assume the encoding state to be pure, $\eta_{I}(\mathbf{z})=|\psi(\mathbf{z})\rangle\langle\psi(\mathbf{z})|$, with the state vector $|\psi(\mathbf{z})\rangle\in\mathcal{H}$ given by
@@
in terms of the encoding unitary $S(\mathbf{z})$ and the initial state $|\Phi_{0}\rangle=|0\rangle^{\otimes N}$. The reservoir unitary $U$ is defined by gate operations or Hamiltonian time evolution, e.g., $U(t)=\exp(-\mathrm{i}tH_{R})$ with a reservoir Hamiltonian $H_{R}$. Specifically, we consider a transverse field Ising Hamiltonian with coupling parameters $h$ and $J$:
@@ III.2 Encoding schemes
This implies a useful data scale of $z\in[0,1]$, since $z\in[1,2]$ produces only an irrelevant global phase relative to $z\in[0,1]$. An $R_{X}$ gate acts similarly, but introduces a complex phase, which is undesired and not needed for classical data processing. $R_{Z}$ gates alone are not suitable for encoding as they just shift the phase of the $|0\rangle$ state. They can, however, be meaningfully combined with $R_{Y}$ gates to encode two scalars into the same qubit:
$\widehat{S}_{j}(z_{j},z_{j+1})=R_{Y}(2\pi z_{j})R_{Z}(2\pi z_{j+1}),$ (27)
@@
$|\psi_{\text{amp}}(z)\rangle=\sqrt{z}|0\rangle+\sqrt{1-z}|1\rangle$ (28)
@@ III.3 Measurement readout
$f^{(c)}(\mathbf{x})=\mathbf{W}^{\top}\mathbf{R}(\mathbf{x}),\qquad\mathbf{W}=[\mathbf{w}^{(1)},\ldots,\mathbf{w}^{(C)}]\in\mathbb{R}^{m_{\text{ro}}\times C}.$ (33)
@@ IV Numerical experiments
In the following, we present numerical experiments on QRCs and QELMs trained using the kernel-based measurement operator optimization. We will typically show test accuracy as a function of the number of measurement operators kept after thresholding. The maximum number of operators resulting from the diagonalization and Pauli projection methods [Sections II.3.2 and II.3.3] is $m_{\text{diag}}=2^{N}\times D_{\text{out}}$ and $m_{\text{pauli}}=4^{N}\times D_{\text{out}}$, respectively, where $D_{\text{out}}$ is the output dimension. In the case of setups without reservoir unitaries, the effective maximum number is lower, since expectation values involving Pauli-$Y$’s vanish identically (see Appendix G). After having determined the optimal operator set, we generally re-fit the QRC using this set [alternatively, one may use Eq. 12].

### v2

@@ Abstract
Finding optimal measurement operators is crucial for the performance of quantum reservoir computers (QRCs), since they employ a fixed quantum feature map. We formulate the training of both stateless (quantum extreme learning machines, QELMs) and stateful (memory dependent) QRCs in the framework of kernel ridge regression. We thus extend the kernel viewpoint of supervised quantum models to recurrent QRCs by deriving an exact Hilbert–Schmidt kernel representation of the optimal readout observable on history space. This approach renders an optimal measurement operator that minimizes prediction error for a given reservoir and training dataset. For large qubit numbers, this method is more efficient than the conventional training of QRCs. We discuss efficiency and practical implementation strategies, including Pauli basis decomposition and operator diagonalization, to adapt the optimal observable to hardware constraints. To demonstrate the effectiveness of this approach, we present numerical experiments on image classification and time series prediction tasks, including chaotic and strongly non-Markovian systems. The developed method can also be applied to other quantum machine learning models.
@@ I Introduction
Since quantum systems are inherently disturbed by measurements, finding efficient and optimal ways to extract information is a crucial part of any quantum machine learning (QML) model, and numerous studies have addressed this problem [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. For variational QML models, measurement operators are often chosen based on hardware constraints or heuristics. Due to the training of the internal layers, this usually does not pose a severe problem for the accuracy of the model. However, in the case of QRCs, training happens only at the readout and thus depends crucially on the information extracted by the measurement operator. Finding optimal measurement operators is thus a key challenge in QRC design [15, 16, 17, 18, 19].
@@
In the present work, we discuss the problem of determining an optimal measurement operator for QRCs using kernel-based methods. QRCs have been introduced as a hardware efficient and noise-resilient quantum ML approach [32, 33, 34, 35], and have been successfully applied in tasks such as image classification [36, 37] and time series prediction [32, 16, 38]. We demonstrate that using an optimal measurement operator enables systematic improvements of QRC performance in these tasks. Besides considering QELMs, which employ a memoryless quantum feature map and are thus close to untrained conventional quantum ML models  [^1] By “conventional”, we mean quantum ML models that have trainable internal layers instead of only a final trainable classical regression layer., we show that our method can also be applied to recurrent QRCs with internal memory. The latter systems are typically used for time series processing [32, 33] and we show that the kernel framework underlying supervised quantum models naturally extends to the recurrent setting by constructing an exact Hilbert–Schmidt kernel on history space.
@@ II.1 Model
We consider an $N$-qubit model, with dimension $D=2^{N}$ of its Hilbert space $\mathcal{H}$, described by the feature map
@@
$f(\mathbf{x})=\sum_{k=1}^{m}w_{k}\mathrm{tr}[M_{k}\rho(\mathbf{x})],$ (2)
@@
Importantly, for the present purpose, the feature map (1) is assumed to be stateless in the sense that the state $\rho$ is always fully determined by the input vector $\mathbf{x}$. Such a model is naturally realized by a quantum extreme learning machine (QELM) [33, 34, 36]. This seems at first to preclude QRCs with internal memory, as commonly used for time series processing [32, 33, 34], because their current state $\rho(t)$ would not only depend on the present input, but also on the inputs preceding it. To address this issue, let $\{\mathbf{z}_{0},\mathbf{z}_{1},\ldots,\mathbf{z}_{T}\}$ be a time series ($T+1$ samples in $\mathbb{R}^{d^{\prime}}$) and assume that the QRC has a memory of length $L$. This approximation is typically warranted due to the echo state property of a QRC. Then, an effective input $x_{t}$ at time $t\geq L$ can be defined as
@@
This quantity can be understood as an ordered set of $L+1$ samples $\mathbf{z}_{t-L},\ldots,\mathbf{z}_{t}$ such that processing these samples through the QRC is sufficient to obtain the unique state $\rho(t)$ associated with $x_{t}$. Accordingly, the QRC output is fully determined by any of the P=TLP=T-L unordered effective inputs $\{x_{t}\}_{t=L}^{T-1}$ and can thus be considered stateless in this setting  [^2] In practice, an initial washout/equilibration time for the QRC must be considered, which, however, does not affect the present theoretical framework.. Prescriptions like (3) to define truncated input histories are also used in kernelization of classical reservoir computing models [30, 31].
@@ II.2 Measurement operator optimization
Trainable object $\mathbf{w}\in\mathbb{R}^{m}$ $M\in\mathcal{M}(\mathcal{H})$ $\boldsymbol{\alpha}\in\mathbb{R}^{P}$
@@
Table 2: Three training formulations for a QRC. The training objective and solution are specialized here to the case of least-squares loss and assuming an orthonormal operator basis. For other kinds of losses, usually no analytic expressions for the weights $\mathbf{w}^{*}$ exist. If the QRC carries an internal memory state, the above formulations remain intact but the data $\mathbf{x}_{i}\hat{\equiv}x_{t}$ represents an input history [see Eq. 3] and $\rho(\mathbf{x})\hat{\equiv}\rho(x_{t})$ is the state resulting by sequentially processing that history (see Sections III.1 and F).
@@
with the same loss function $\mathcal{L}$ as in Eq. 4. In practice, this optimization over $\mathcal{M}(\mathcal{H})$ can be achieved by using in (4) as observables $\{M_{k}\}$ a complete orthonormal $m=D^{2}$-dimensional basis of $\mathcal{M}(\mathcal{H})$ (e.g., the normalized Pauli basis $\{2^{-N/2}P_{k}\}$). The associated optimal weights $w_{k}^{*}$ are given formally by the same expression as in Eq. 5, and yield the optimal observable
@@
The above unconstrained optimization procedure becomes especially natural in its dual formulation, resting on the connection between supervised quantum ML models and quantum kernels [20]. Identifying the unconstrained feature map as $\phi(\mathbf{x})=\rho(\mathbf{x})$ allows one to define the kernel $K(\mathbf{x},\mathbf{x}^{\prime})=\mathrm{tr}(\rho(\mathbf{x})\rho(\mathbf{x}^{\prime}))$, which induces a reproducing kernel Hilbert space (RKHS) [42]. The representer theorem implies that the optimizer $M^{*}$ (optimal measurement operator) of the problem (6) is of the form (see also Appendix A)
$M^{*}=\sum_{j=1}^{P}\alpha_{j}^{*}\rho(\mathbf{x}_{j}),$ (8)
@@
For the least-squares loss, the canonical solution coefficients are available in closed form as
$\boldsymbol{\alpha}^{*}=(\mathbf{K}+\lambda\mathbbm{1})^{-1}\mathbf{y},\qquad K_{ij}=\mathrm{tr}(\rho(\mathbf{x}_{i})\rho(\mathbf{x}_{j})).$ (10)
For other loss functions, the solution coefficients must typically be computed numerically. Independently of the loss function $\mathcal{L}$, an optimally trained QELM [Eq. 2] admits the kernel form
@@
The generalization to multi-dimensional target values $\mathbf{y}_{i}\in\mathbb{R}^{C}$ ($i=1,\ldots,P$) is straightforward and requires a separate kernel construction like in Eq. 11 for each class (this includes multi-class classification in a one-vs-rest scheme).
When the measurement operators $\{M_{k}\}$ form a complete orthonormal basis (or at least span the relevant subspace containing $M^{*}$ [17]), primal and dual formulations are equivalent and Eqs. 7 and 8 describe the same object. Expanding the optimal observable $M^{*}$ in a complete orthogonal basis leads to a relation between the coefficients $\mathbf{w}^{*}$ and $\boldsymbol{\alpha}^{*}$ via [cf. Eqs. 49 and 61]
$w_{k}^{*}=\frac{1}{g_{k}}\sum_{j=1}^{P}\alpha_{j}^{*}\mathrm{tr}(M_{k}\rho(\mathbf{x}_{j})).$ (12)
Here, $\mathrm{tr}(M_{k}M_{l})=g_{k}\delta_{kl}$ with $g_{k}=2^{N}$ for the orthogonal Pauli basis. The optimization formulations are summarized in Table 2.
@@
(i) Neither the primal nor the dual optimization methods assume uncorrelated samples. Both are therefore compatible with the sliding window construction (3). (ii) Global conjugation $\rho(\mathbf{x})\mapsto U\,\rho(\mathbf{x})\,U^{\dagger}$ leaves the expression (11) unchanged and only transforms the optimal observable as $M^{*}\mapsto UM^{*}U^{\dagger}$. (iii) The constrained primal formulation can be “dualized”, too, by using the kernel induced by the constrained feature map (see Table 2). While this does not render a lower training error, it may offer computational benefits.
@@ II.3.3 Diagonalization
The prediction is the weighted average of the observed bitstrings using the eigenvalues $\{\lambda_{\mathbf{b}}\}$. Thus, the optimal observable can be implemented by a single basis rotation $V^{\dagger}$ followed by a standard computational-basis measurement. This is equivalent to measuring all $2^{N}$ distinct Pauli-$Z$ strings $\{Z_{\mathbf{s}}\}_{\mathbf{s}\in\{0,1\}^{N}}$, since (see Appendix C)
@@
The unitary $V$ may be treated as part of the readout (applied once just before measurement) or absorbed into the reservoir by redefining the last layer as $U_{R}\mapsto V^{\dagger}U_{R}$. This global conjugation leaves the kernel $K(\mathbf{x},\mathbf{x}^{\prime})$ and the trained predictor unchanged (see Section II.2) and only affects the hardware realization of the optimal measurement. While, strictly, a pre-rotation followed by computational-basis measurement is sufficient and equivalent to measuring the exact $M^{*}$, we will include in our numerical experiments [Section IV] also the Pauli observable decomposition [Eq. 13], as both approaches can lead to slight differences in numerical stability.
@@ III.1 QRCs and QELMs
The standard (stateful) QRC model introduced in [32, 33] consists of a set of coherently evolving (memory/ancilla) qubits ($N_{A}$) periodically coupled to a set of input qubits ($N_{I}$). At time $t$, a classical input $\mathbf{z}_{t}\in\mathbb{R}^{d^{\prime}}$ is encoded into an input state with density matrix $\eta_{I}(\mathbf{z}_{t})$. The joint state with the coherent ancilla state $\eta_{M}(t)$ is formed and time-evolved by a unitary $U$:
$\eta(t)=U\left(\eta_{I}(\mathbf{z}_{t})\otimes\eta_{M}(t)\right)U^{\dagger}.$ (17)
This $\eta(t)$ represents the feature map $\rho(x_{t})$ on the input sequence defined by Eq. 3 [corresponding to the one in (1)]. Subsequently, the input is reset and discarded, and the memory alone is carried to the next step:
$\eta_{M}(t+1)=\mathrm{tr}_{I}\!\left[\eta(t)\right],$ (18)
where $\mathrm{tr}_{I}$ denotes partial trace over the input register $I$ (see Appendix F for further details).
We assume the encoding state to be pure, $\eta_{I}(\mathbf{z})=|\psi(\mathbf{z})\rangle\langle\psi(\mathbf{z})|$, with the state vector $|\psi(\mathbf{z})\rangle\in\mathcal{H}_{I}$ (living in the $N_{I}$-qubit input Hilbert space) given by
@@
in terms of the encoding unitary $S(\mathbf{z})$ and the initial state $|\Phi_{0}\rangle=|0\rangle^{\otimes N_{I}}$. The reservoir unitary $U$ is defined by gate operations or Hamiltonian time evolution, e.g., $U(t)=\exp(-\mathrm{i}tH_{R})$ with a reservoir Hamiltonian $H_{R}$. Specifically, we consider a transverse field Ising Hamiltonian with coupling parameters $h$ and $J$:
@@ III.2 Encoding schemes
This implies a useful data scale of $z\in[0,1]$, since $z\in[1,2]$ produces only an irrelevant global phase relative to $z\in[0,1]$. An $R_{X}$ gate acts similarly, but introduces a complex phase, which is undesired and not needed for classical data processing. $R_{Z}$ gates alone are not suitable for encoding as they just shift the phase of the $|0\rangle$ state. They can, however, be meaningfully combined with $R_{Y}$ gates to encode two scalars into the same qubit (dense encoding):
$\widehat{S}_{j}(z_{j},z_{j+1})=R_{Z}(2\pi z_{j+1})R_{Y}(2\pi z_{j}),$ (27)
@@
$|\psi_{\text{amp}}(z)\rangle=\sqrt{z}|0\rangle+\sqrt{1-z}|1\rangle.$ (28)
@@ III.3 Measurement readout
$\mathbf{f}(\mathbf{x})=\mathbf{W}^{\top}\mathbf{R}(\mathbf{x}),\qquad\mathbf{W}=[\mathbf{w}^{(1)},\ldots,\mathbf{w}^{(C)}]\in\mathbb{R}^{m_{\text{ro}}\times C}.$ (33)
Note that the measurement-operator optimization theory developed here applies only to the linear readout case.
@@ IV Numerical experiments
In the following, we present numerical experiments on QRCs and QELMs trained using the kernel-based measurement operator optimization. We will typically show test accuracy as a function of the number of measurement operators kept after thresholding. The maximum number of operators resulting from the diagonalization and Pauli projection methods [Sections II.3.2 and II.3.3] is $m_{\text{diag}}=2^{N}\times D_{\text{out}}$ and $m_{\text{pauli}}=4^{N}\times D_{\text{out}}$, respectively, where $D_{\text{out}}$ is the output dimension. In the case of setups without reservoir unitaries, the effective maximum number can be lower, since expectation values involving Pauli-$Y$’s vanish identically on states encoded via (26), (28), or (29) (see Appendix G). After having determined the optimal operator set, we generally re-fit the QRC using this set [alternatively, one may use Eq. 12].